OpenStudy (liliakarina):

How to simplify sqrt24/300 ?

OpenStudy (liliakarina):

\[\sqrt{24/300}\]

mathslover (mathslover):

\[\huge{\sqrt{\frac{24}{300}}}\] \[\huge{\sqrt{\frac{4*6}{100*3}}}\] \[\huge{\frac{2}{10}\sqrt{\frac{6}{3}}}\] \[\huge{\frac{2}{10}\sqrt{2}}\]

OpenStudy (liliakarina):

Could it be the equivalent to \[5/\sqrt{2}\]

mathslover (mathslover):

Sorry it can be : \[\huge{\frac{\sqrt{2}}{5}}\]

mathslover (mathslover):